OpenStudy (anonymous):

An isosceles triangle has an area of 156ft^2. If the base is 16, what are the lengths of the legs?

OpenStudy (anonymous):

Area = ½ × b × h b = base h = vertical height

OpenStudy (anonymous):



OpenStudy (anonymous):

Area = ½ × b × h 156 = 1/2 * 16 * h 156 = 8 * h ---- ----- 8 8 19.5 = h

OpenStudy (anonymous):

so then both legs equal 19.5 right?or is that the height...